OpenStudy (anonymous):

how do you slove this? 3d+4=2+3d-1/2 is the answer no solution?

jimthompson5910 (jim_thompson5910):

if the equation is \[\large 3d + 4 = 2 + 3d - \frac{1}{2}\] then you are correct in saying "no solutions"

OpenStudy (jdoe0001):

\(\bf \cancel{3d} + 4 = 2 + \cancel{3d} - \cfrac{1}{2} \implies 4 = 2 - \cfrac{1}{2} \implies 4 \ne \cfrac{3}{2}\)